408juju:

Given the function h(x) = 3(5)x, Section A is from x = 0 to x = 1 and Section B is from x = 2 to x = 3. Part A: Find the average rate of change of each section. (4 points) Part B: How many times greater is the average rate of change of Section B than Section A? Explain why one rate of change is greater than the other. (6 points) if you help me with this ill give you modded cars on gta ps4 :)

jimthompson5910:

Is that x an exponent?

408juju:

one sec

408juju:

which x ?

408juju:

h(x) ?

jimthompson5910:

when you wrote 3(5)x

408juju:

i dont think so

jimthompson5910:

I'm curious if it says \[3(5)^x \] the x being higher than the 5 and smaller in font size too

408juju:

its not like that on the question its above it if thats what you mean

jimthompson5910:

If you can please post a screenshot of the full problem

408juju:

ok lemme find out how to screen shot rq im on school computer

408juju:

jimthompson5910:

Thanks that clears up everything

408juju:

your welcome dad

408juju:

sir*

jimthompson5910:

So what we need to do first is plug in x = 1 into h(x) What do you get when you do that?

408juju:

how would you plug it in ?

408juju:

dont give answer just explain how to

jimthompson5910:

you replace every x with 1 Then use PEMDAS or your calculator to evaluate h(x) = 3(5)^x h(1) = 3(5)^1 h(1) = ??

408juju:

one second please

jimthompson5910:

sure thing

408juju:

im getting help on this lmao

408juju:

45?

jimthompson5910:

what is 5^1 equal to

jimthompson5910:

5^1 means 5 to the power of 1 We can write that as \(5^1\)

408juju:

the answer is 45

jimthompson5910:

I'm not sure how you're getting 45. Is your calculator saying \[5^1 = 45\] or something else?

408juju:

wdyem

408juju:

wdyem= what do you even mean

408juju:

i got the answer but i just need to know if its right

408juju:

you there????

408juju:

did you leave to get the milk again???? :(